Find the future value, using the future value formula and a calculator. $889 at 5.5% compounded quarterly for 5 years
Answer by TimothyLamb(4379)

compound-interest formula:
v = p( 1 + (r/n) )^(nt)
where:
v = final value
p = initial (principal) value
r = interest rate (as a decimal)
n = number of compounding periods per year
t = number of years
---
v = 889*( 1 + (0.055/4) )^(4*5)
v = 889*( 1 + (0.055/4) )^20
v = 1168.21
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---